OpenStudy (exoexoexo):

if the point (3,-7) is the center of a circle and the point (8,5) is on the circle, what is the circumference of the circle?

OpenStudy (jdoe0001):

 find the radius keeping in mind that \(\bf \textit{circumference}=2\pi r\qquad r=radius\)

OpenStudy (jdoe0001):

so the radius is really the distance between, the center point, and point on the circle thus \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 3}}\quad ,&{\color{blue}{ -7}})\quad % (c,d) &({\color{red}{ 8}}\quad ,&{\color{blue}{ 5}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\to radius\)